OpenStudy (anonymous):

3y+2x=sin(xy^2) Find dy/dx Can someone help me please

OpenStudy (anonymous):

theres a product rule inside the sin(xy^2)

OpenStudy (anonymous):

Implicit differentiation

OpenStudy (anonymous):

yeah it comes out wrong i dont know why

OpenStudy (anonymous):

left side i get 3yy'+2

OpenStudy (anonymous):

and the right side i get cosy^2+cosy^2y'x

OpenStudy (anonymous):

whats sham ?

OpenStudy (shamil98):

i already know how to do this, it's calc 1 bat -.-

OpenStudy (shamil98):

let's see you walk her through it. im busy atm

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

i get other questions similar to this right this one i keep getting it wrong

OpenStudy (anonymous):

so the right side is 3yy'+2=cosy^2+cos2yy'x

OpenStudy (anonymous):

than i combine like terms?

OpenStudy (anonymous):

\[\frac{ d }{ dx }(2x+3y) = \frac{ d }{ dx }(\sin(xy^2))\] Alright I'm going to show you this step by step, so you don't make a mistake next time you do a question like this, do you like the d/dx notation or do you want me to use y'?

OpenStudy (anonymous):

use y' and thanx :)

OpenStudy (anonymous):

\[(2x+3y)'=(\sin(xy^2))'\] I just saw you mention y' so I deleted everything I did lol, I'll restart though this way, it's fine. Haha.

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

\[2+3y'=\cos(xy^2)*(xy^2)'\] chain rule \[2+3y' = \cos(xy^2) *((x)'(y^2)+(y^2)'(x))\] \[2+3y' = \cos(xy^2)*(y^2+2yy'x)\] \[2+3y' = \cos(xy^2)y^2+2xcos(xy^2)yy'\] Expand \[2+3y'-2xcos(xy^2)yy'=\cos(xy^2)y^2\] Subtracting \[2xyy'\cos(xy^2)\] from sides \[3y'-2xcos(xy^2)yy'=-2+\cos(xy^2)y^2\] Subtracting 2. \[y'=\frac{ -2+\cos(xy^2)y^2 }{ 3-2xcos(xy^2)y }\] Dividing both sides by \[3-2xycos(xy^2)\]

OpenStudy (anonymous):

Hope that helps :p

OpenStudy (anonymous):

Thanx!

OpenStudy (anonymous):

Your welcome :)